OpenStudy (anonymous):

PR has an endpoint at (25, -5) and a midpoint of (18, -1). What is the value of the x-coordinate of the other endpoint?

OpenStudy (lgbasallote):

the x coordinate of the midpoint is equal to \[\Large \frac{x_1 + x_2}{2}\] let your other endpont be equal to x and the given endpoint be equal to 25 \[\Large \frac{x+25}{2} = 18\] agree?

OpenStudy (anonymous):

so how do i solve it ?